OpenStudy (anonymous):

Rewrite as a logarithmic equation: 2⁴=(a-b)

OpenStudy (anonymous):

\[\log_{2}(a-b)=4\]

OpenStudy (mathstudent55):

\[y = b ^{x} \space <------> \space \log _{b}y = x\] The base of a power, b, is the base of the log. What the log equals, x, is the exponent. What the power equals, y, is what you're taking the log of. So here you get: \[\log _{2} (a - b) = 4\]

OpenStudy (anonymous):

@mathstudent55 i understand much better now, thank you so much!!

OpenStudy (mathstudent55):

Keep in mind two things, the base of the power is the base of the log. The log of the quantity is the exponent you need to raise the base to.

OpenStudy (mathstudent55):

You're welcome